. Solve the following LP minimization problem.

        Min    3X + 2Y
        s.t.    5X + 3Y <= 30
                 3X + 4Y >= 36
                   Y >= 7
             X , Y >= 0

Group of answer choices

X = 0, Y= 9

The optimal value of the objective function is 5.

None of the other answers are correct.

The optimal value of the objective function is 7.

X = 1, Y = 3
Answer #1

min 3x + 2y 5X +34 < 30 3x + 4 V 236 Y 27 X, V 20 Constraint o 5x+34 L30 at Y=o, a point is 5x + 0 = 30 ..(x=6) so (60) is afor constraint ② points are (12,0) & (0,9) th the feasibu vagen * the feasible region is B solving constraint 0 A = (0,92, B=Volf C-5(1) Y = 8.18 8.18 Points objective function 7= 3x +24 minimum value is to be determined at point A=(0,9) Z = 310 + 2(min { le, 20, 19.63) minimum objective = 18 - value at FA at a points are (0,9). at points of (x=0 and V=9 Construction Con